An unstable atomic nucleus with with a mass of 1.7x10^-26kg is at rest at the origin of a coordinate system. It spontaneously disintegrates into three particles: one moving in the positive y-direction with a mass of 5x10^-27 kg and a speed of 6x10^6 m/s, the second moves in the positive x-direction with a mass of 8.4x10^-27kg and a speed of 4x10^6m/s. Find the magnitude and direction of the velocity of the third particle.
Answer #1

Solution: mass of third particle is 1 m₂ = M-(m, tma) = 1-7X1026 (3x10277 8.481027) = 3.6810-27 kg Since, the nuclers is at r